# Marvel: Accelerating Safe Online Reinforcement Learning with Finetuned Offline Policy

### v2

@@ 4 Experiments
In this section, we conduct extensive experiments to verify the effectiveness of our approach, aiming to answer the following questions: 1) RQ1: How does our method compare with naive finetuning and other SOTA baselines in both reward and cost? 2) RQ2: How do different components in Marvel affect the performance? Due to the space limit, we delegate the experimental details and some additional results to Appendix D and Appendix F.
@@ 4.1 Evaluation Setup
Benchmarks. We consider the DSRL benchmark [10] and select ten environments from the Bullet Safety Gym [17] and Safety Gymnasium [33]: BallRun, BallCircle, CarRun, CarCircle, HalfCheetah, AntCircle, AntRun, DroneCircle, Hopper, and Swimmer (results for the last four are in Section D.1). The cost threshold is set to be 20 in these environments. As mentioned earlier in Section 2, we choose CPQ and SAC-lag as base algorithms in our proposed framework Marvel for offline training and online finetuning, respectively, due to the effectiveness and representativeness of them. Each experiment was conducted using five random seeds, and the results were averaged to generate the final learning curves. We use a dataset that includes data provided by DSRL [34] and random data generated by a random policy to control the quality of the offline dataset.
@@
Baselines. While Guided Online Distillation [25] is the only work studying O2O safe RL, its usage of large pretrained model leads to an unfair comparison with standard RL frameworks using typically small-scale policy networks. In this work, we compare Marvel with JSRL [27], as Guided Online Distillation mainly follows this approach except using DT as the pretrained policy. Besides, we further adapt some SOTA approaches in O2O RL to O2O safe RL, including SO2 [35] and PEX [36], and a Warm Start approach as baselines. SO2 improves Q-value estimation through Perturbed Value Updates, JSRL and PEX utilize offline pretrained policies for exploration, and Warm Start directly finetunes the policy and Q-networks from offline safe RL without modifications. We also compare with online learning from scratch, namely From Scratch. More importantly, aPID is used to update Lagrange multipliers in all these baseline methods, which in fact already improves the performance of these methods compared to their original designs. These baselines provide a meaningful comparison to demonstrate the effectiveness of Marvel in O2O safe RL. We provide the detailed descriptions of each baseline algorithm in Appendix C.
@@ 4.2 Main Results
As shown in Fig. 3, Marvel demonstrates better or comparable performance compared to all baselines consistently across all environments, i.e., achieving the higher return while keeping the cost below the threshold. In stark contrast, the naive warm start method proves largely ineffective, often causing performance drop or stagnation during training. Without aligning the Q-estimations, both JSRL and PEX struggles a lot to improve during online learning and fails to control the cost. Besides, PEX also suffers from poor training stability and high variance across different settings. While SO2 mitigates the inaccuracies of Q-estimations related to O2O RL, it does so only to a limited extent and cannot maintain its performance consistently across different environments, although aPID has already been used to boost its performance. On the other hand, the fact that SO2 performs better than other baselines further indicates 1) the great potentials of enabling fast and safe online learning through policy finetuning (compared to using the pretrained policy only as a guide policy as in JSRL and PEX) and 2) the need of correcting pretrained Q-estimations before online finetuning.
More importantly, Marvel shows the superior capability of finding a good and safe online policy very quickly by using only a few steps of online interactions. In particularly, in environments like BallCircle and CarCircle, Marvel finds a good policy within less than $15$ steps, dominating baseline methods in both performance and speed. Note that all approaches indeed start from the offline policy and Q-functions, i.e., the same point at step 0 (ignored in all figures). Guided by the offline policy and aligned Q-functions, Marvel rapidly jumps into the high-reward region in the state space, which highlights the effectiveness of VPA in addressing the overly conservative nature of offline pretrained policies. But this may also lead to a high cost at the beginning, e.g., the cost spike in the early step of finetuning as illustrated in Fig. 3. Because a few steps of online finetuning will just modify the policy in a neighborhood of the pretrained policy, aPID will lead the finetuned policy to low cost state-actions in the high-reward region by adaptively pushing the cost towards the threshold. It is also worth to note that we use the same aPID parameters across all tested environments without any further adjustments, demonstrating the robustness and effectiveness of our design.
Figure 4: The policy pretrained with BEAR-lag performs poorly in cost, as this algorithm was not designed for safe RL. However, Marvel still achieves good results. This also indicates that Marvel performs well with pretrained policies of varying initial performance.
@@
Compatibility of Marvel: In O2O safe RL, compatibility with different offline safe RL methods is essential. Given the non-interactive nature of offline training and the potential unavailability of algorithms due to privacy concerns, this compatibility becomes even more critical compared to online algorithms. Our design of Marvel naturally fits a variety of offline safe RL methods and only requires a pretrained policy and Q-functions. To further verify this, Fig. 4 shows the training process using BEAR+Lagrangian (BEAR-lag) [37] in the offline phase and SAC-lag in online finetuning. Note that BEAR-lag was not specifically designed for offline safe RL, but rather it incorporates the Lagrange multiplier into offline RL. In the BallCircle environment, Marvel achieves the highest reward while satisfying cost constraints, and in the CarRun environment, it also outperformed others while maintaining cost below the threshold. This highlights the flexibility of our algorithm across different offline safe RL methods.
@@ 4.3 Ablation Studies
To answer RQ2, we conduct experiments in various setups. As shown in Fig. 5, the performance is best when both VPA and aPID are used. In contrast, if only VPA is used with traditional dual ascent during online finetuning, it significantly slows down safe online learning and takes a much longer time to reduce the cost. If only aPID is applied without VPA, the learning performance is very similar to naive policy finetuning, which struggles to improve due to the erroneous Q-estimations. We also evaluate the effectiveness of adaptive control in aPID, by comparing the performance between Marvel (VPA+aPID) and Marvel with aPID replaced by PID (VPA+PID). It can be seen from Fig. 5 that the training curve for cost exhibits significant fluctuations without using aPID. More critically, when the cost is close to the limit, PID cannot reduce its control strength. As a result, even if on average the cost of VPA+PID is close to the threshold, it is very frequent that the real-time cost exceeds the limit substantially, which is in fact not safe. Moreover, the inability to adjust the Lagrange multiplier promptly and appropriately affects the weight of reward and cost in policy updates, thereby influencing reward performance, as shown in the plots for CarCircle and HalfCheetah.
Figure 5: ‘aPID’ represents the scenario where aPID is used exclusively during the online finetuning phase, without applying VPA. ‘VPA’, on the other hand, shows the case where aPID is not used, and VPA is applied only during the pre-finetuning phase. VPA+PID use PID control for finetuning, while VPA+aPID employs aPID with adaptive parameter adjustment. Clearly, VPA+aPID achieves the best performance in terms of learning performance, speed, and stability.
We also provide additional ablation studies in the Section D.2, which can provide the following insights: 1) Applying VPA to both the pretrained reward and cost Q-functions achieves the best performance compared to applying VPA to only one of them, which is reasonable as both will be manipulated during offline learning to reduce the extrapolation errors. 2) Marvel is robust to different qualities of the offline dataset. Regardless of the performance of the offline pretrained policy, Marvel can effectively finetune the policy while keeping the cost below the limit. 3) Introducing the entropy term in VPA is very helpful to improve the policy performance in terms of reward, by encouraging explorations of high-entropy states.
@@ 5 Conclusion
O2O safe RL has great potentials to put safe RL on the ground in real-world applications, by leveraging offline learning to facilitate fast online safe learning. In this paper, we proposed the first policy-finetuning based framework, namely Marvel, for O2O safe RL. In particular, by showing that naive finetuning would not work well, we identified two unique challenges in O2O safe RL, i.e., the erroneous Q-estimations and Lagrangian mismatch. To address these challenges, Marvel consisted of two key designs: 1) value pre-alignment to correct the Q-estimations before online finetuning, and 2) adaptive PID control to dynamically change the control parameters so as to rapidly and appropriately control the cost. Extensive experiments demonstrate the superiority of Marvel over multiple baselines. More importantly, Marvel is compatible to a variety of offline and online safe RL approaches, making it very practically appealing. For future work, it is interesting to take a closer look at the offline dataset, to identify states that are more worth exploring during online finetuning given the environmental information and cost threshold. Ultimately, we hope our work will bridge the gap between offline and online algorithms in safe RL, distinct from unconstrained RL, and enhance the efficiency of online safe RL, laying the foundation for the usage of safe RL in practical applications.
@@ Appendix A Overview of Algorithm
Data: Offline RL algorithm $\{L_{off}^{Q_{\phi}},L_{off}^{Q_{c_{\phi_{c}}}},L_{off}^{\pi_{\theta}}\}$, Online RL algorithm $\{L_{on}^{Q_{\phi}},L_{on}^{Q_{c_{\phi_{c}}}},L_{on}^{\pi_{\theta}}\}$
Result: Offline dataset $\mathcal{D}_{off}$, Online dataset $\mathcal{D}_{on}$, Network parameters $\phi,\phi_{c},\theta$, Lagrange multiplier $\lambda$
while in offline training phase do
       $\phi\leftarrow\phi-\lambda_{Q}\nabla_{\phi}L_{off}^{Q_{\phi}}$;
       $\phi_{c}\leftarrow\phi_{c}-\lambda_{Q}^{c}\nabla_{\phi_{c}}L_{off}^{Q_{c_{\phi_{c}}}}$;
       $\theta\leftarrow\theta-\lambda_{\pi}\nabla_{\theta}L_{off}^{\pi_{\theta}}$;
end while
% VPA with offline dataset;
while in VPA phase do
       for each VPA step do
             Sample transitions $(s,a,r,c,s^{\prime})\sim\mathcal{D}_{off}$;
             Update $Q$ by Eq. 8, Update $Q_{c}$ by Eq. 9;
       end for
end while
while in online training phase do
       for each environment step do
             $a_{t}\sim\pi_{\theta}(s_{t})$, $s_{t+1}\sim T(s_{t},a_{t})$;
             $\mathcal{D}_{on}=\mathcal{D}_{on}\cup\{s_{t},a_{t},r(s_{t},a_{t}),c(s_{t},a_{t}),s_{t+1}\}$;
       end for
      for each update step do
             $\phi\leftarrow\phi-\lambda_{Q}\nabla_{\phi}L_{on}^{Q_{\phi}}$;
             $\phi_{c}\leftarrow\phi_{c}-\lambda_{Q}^{c}\nabla_{\phi_{c}}L_{on}^{Q_{c_{\phi_{c}}}}$;
             $\theta\leftarrow\theta-\lambda_{\pi}\nabla_{\theta}L_{on}^{\pi_{\theta}}$;
             % update Lagrange multiplier with aPID;
             Update $\lambda$ by Eq. 10;
             update PID parameters by Eq. 11 and Eq. 12;
       end for
end while

### v3

@@ 4 Experiments
The objective of O2O safe RL is to utilize offline data to enable fast online finetuning, i.e., maximizing rewards while satisfying the cost threshold, with minimal interactions with the environment. Therefore, an important evaluation criterion for effective O2O safe RL methods is whether the method can quickly obtain a good policy with high rewards and low safety violations only after a limited number of online interactions, which is the setup considered in this paper. Due to the space limit, we delegate the experimental details and some additional results to Appendix D and Appendix G.
@@ 4.1 Evaluation Setup
Benchmarks. We consider the DSRL benchmark (Liu et al., 2023b) and select ten environments from the Bullet Safety Gym (Gronauer, 2022) and Safety Gymnasium (Ji et al., 2023): BallRun, BallCircle, CarRun, CarCircle, HalfCheetah, AntCircle, AntRun, DroneCircle, Hopper, and Swimmer (results for the last four are in Appendix D.1). The cost threshold is set to be 20 in these environments. As mentioned earlier in Section 2, we choose CPQ and SAC-lag as base algorithms in our proposed framework Marvel for offline training and online finetuning, respectively, due to the effectiveness and representativeness of them. Each experiment was conducted using five random seeds, and the results were averaged to generate the final learning curves. We use a dataset that includes data provided by DSRL (Liu et al., 2024) and random data generated by a random policy to control the quality of the offline dataset.
@@
Baselines. While Guided Online Distillation (Li et al., 2024) is the only work studying O2O safe RL, its usage of large pretrained model leads to an unfair comparison with standard RL frameworks using typically small-scale policy networks. In this work, we compare Marvel with JSRL (Uchendu et al., 2023), as Guided Online Distillation mainly follows this approach except using DT as the pretrained policy. Besides, we further adapt some SOTA approaches in O2O RL to O2O safe RL, including SO2 (Zhang et al., 2024) and PEX (Zhang et al., 2023a), and a Warm Start approach as baselines. SO2 improves Q-value estimation through Perturbed Value Updates, JSRL and PEX utilize offline pretrained policies for exploration, and Warm Start directly finetunes the policy and Q-networks from offline safe RL without modifications. We compare with online learning from scratch, namely From Scratch, i.e., SAC-lag, which can achieve strong performance but require a large number of online interactions. For completeness, we also evaluate the performance of Cal-QL (Nakamoto et al., 2023) and SAC-Lag with an offline replay buffer across four representative environments: BallCircle, BallRun, CarCircle, and CarRun. The corresponding results are presented in Fig. 6. Given the interaction and network update settings in our experiments, we compare our implementation of SAC-lag with the standard FSRL library (Liu et al., 2024) in Appendix D.5, and the results show that their performance is highly comparable.
More importantly, aPID is used to update Lagrange multipliers in all these baseline methods (instead of traditional dual ascent), which in fact already improves the performance of these methods compared to their original designs. These baselines provide a meaningful comparison to demonstrate the effectiveness of Marvel, but also indicate the need of addressing the two challenges simultaneously for O2O safe RL. We provide the detailed descriptions of each baseline algorithm in Appendix C. Additionally, we provide a comparison with CDT (Liu et al., 2023c), using a smaller decision transformer (around 600k parameters) than the one in (Li et al., 2024), in Appendix D.1, and also demonstrate Marvel’s superiority over baselines with large models.
Table 2: The “offline" column represents the performance of the offline safe RL pre-training. We present the performance at 20% of the total training steps during the finetuning process, as well as the performance after all the training steps are completed, e.g. “100 steps r/c" and “500 steps r/c" for reward and cost, respectively. Cost values greater than 20 are shown in gray. Bold numbers represent the highest reward while satisfying the cost threshold. The plots are in Appendix D.
Environment Marvel (ours) From Scratch Warm Start SO2 JSRL PEX Offline
BallCircle 100 steps: r/c 580.3 / 19.2 170.8 / 88.9 54.7 / 0.0 54.4 / 0.0 -7.6 / 137.5 0.0 / 119.7 166.0 / 11.0
500 steps: r/c 603.9 / 19.8 241.7 / 10.9 176.6 / 18.5 58.4 / 2.0 1.6 / 165.4 -4.3 / 39.9
BallRun 100 steps: r/c 276.0 / 5.8 1055.9 / 75.4 346.3 / 22.1 338.6 / 19.9 -120.5 / 37.2 270.3 / 42.9 262.0 / 3.0
500 steps: r/c 306.6 / 5.5 315.2 / 5.6 132.2 / 23.0 286.8 / 12.1 174.0 / 92.1 -555.6 / 88.8
CarCircle 100 steps: r/c 330.6 / 18.8 49.1 / 16.4 73.2 / 71.1 69.3 / 17.6 2.4 / 36.3 -0.5 / 69.4 265.0 / 14.0
500 steps: r/c 341.3 / 19.5 115.2 / 12.7 141.6 / 24.1 115.9 / 9.1 1.1 / 130.0 -12.6 / 142.9
CarRun 100 steps: r/c 537.9 / 17.0 337.1 / 74.3 383.9 / 64.4 510.2 / 17.3 -186.4 / 140.8 -183.6 / 112.5 544.0 / 72.0
500 steps: r/c 547.5 / 18.2 293.5 / 7.2 391.4 / 16.9 522.2 / 16.2 126.6 / 128.2 153.0 / 38.8
AntCircle 800 steps: r/c 4.0 / 0.0 1.2 / 0.0 1.1 / 0.0 2.4 / 0.0 0.0 / 0.0 1.3 / 0.2 4.0 / 39.0
4000 steps: r/c 5.6 / 0.9 1.5 / 0.0 1.3 / 0.0 3.7 / 1.3 0.6 / 1.9 6.9 / 2.3
HalfCheetah 800 steps: r/c 1343.6 / 19.2 39.3 / 0.0 1695.9 / 180.5 1169.5 / 27.1 0.1 / 0.0 -70.6 / 0.0 113.0 / 17.0
4000 steps: r/c 1544.8 / 20.0 1074.3 / 28.8 974.1 / 15.4 559.4 / 23.9 27.6 / 3.8 -155.6 / 0.2
BallCircle (BEAR-lag) 100 steps: r/c 621.1 / 19.2 170.8 / 88.9 174.5 / 101.9 -9.8 / 135.1 4.7 / 150.2 24.5 / 129.4 394.0 / 19.0
500 steps: r/c 552.3 / 15.2 241.7 / 10.9 440.2 / 100.0 71.8 / 16.6 10.5 / 163.0 -13.6 / 5.8
CarRun (BEAR-lag) 100 steps: r/c 410.3 / 60.0 337.1 / 74.3 398.7 / 38.1 191.1 / 1.3 125.0 / 84.2 97.8 / 15.2 270.0 / 17.0
500 steps: r/c 510.5 / 19.2 293.5 / 7.2 204.6 / 50.1 -59.6 / 0.0 31.9 / 113.6 200.6 / 34.0
@@ 4.2 Main Results
As shown in Table 2, Marvel demonstrates better or comparable performance compared to all baselines consistently across all environments, i.e., achieving the higher return while keeping the cost below the threshold. In particular, as shown in Fig. 4, Marvel demonstrates superior stability compared to other baseline algorithms. Note that a key aspect of safe RL is ensuring safety, which requires keeping the cost below the predefined threshold while maximizing reward. Marvel achieves this balance effectively, as its cost remains below the threshold in all environments. In stark contrast, the naive warm start method proves largely ineffective, often causing performance drop or stagnation during training. Without aligning the Q-estimations, both JSRL and PEX struggles a lot to improve during online learning and fails to control the cost. While SO2 mitigates the inaccuracies of Q-estimations related to O2O RL, it does so only to a limited extent and cannot maintain its performance consistently across different environments, even though aPID has already been used to boost its performance. On the other hand, the fact that SO2 performs better than other baselines further indicates 1) the great potentials of enabling fast and safe online learning through policy finetuning (compared to using the pretrained policy only as a guide policy as in JSRL and PEX) and 2) the need of correcting pretrained Q-estimations before online finetuning.
More importantly, Marvel demonstrates the ability to find a good and safe online policy very quickly with minimal online interactions, as shown in Fig. 4 and Fig. 5 (shown in Appendix D). For instance, in the BallCircle and CarCircle environments, near-optimal performance can be achieved with merely 20 steps of gradient updates. By leveraging offline policies and aligned Q-functions, Marvel rapidly explores high-reward regions while addressing the overly conservative nature of offline pretrained policies. Although initial finetuning may cause a cost spike (Fig. 4), aPID effectively reduces costs by guiding the policy towards low-cost state-actions in high-reward regions. Also, Marvel achieves the best max reward for a given cumulative cost, illustrated in Fig. 16, demonstrating superior performance while incurring fewer violations in the environment.” Notably, the same aPID parameters are used across all environments, showcasing its robustness and adaptability.
@@
Compatibility of Marvel: In O2O safe RL, compatibility with different offline safe RL methods is essential. Given the non-interactive nature of offline training and the potential unavailability of algorithms due to privacy concerns, this compatibility becomes even more critical compared to online algorithms. Our design of Marvel naturally fits a variety of offline safe RL methods and only requires a pretrained policy and Q-functions. To further verify this, the last two rows of table˜2 show the training process using BEAR+Lagrangian (BEAR-lag) (Kumar et al., 2019) in the offline phase and SAC-lag in online finetuning. Note that BEAR-lag was not specifically designed for offline safe RL, but rather it incorporates the Lagrange multiplier into offline RL. In BallCircle, Marvel achieves the highest reward while satisfying cost constraints, and in the CarRun environment, it also outperformed others while maintaining cost below the threshold. This highlights the flexibility of our algorithm across different offline safe RL methods.
Figure 3: ‘aPID’ refers to adaptive PID applied only during online finetuning; ‘VPA’ refers to VPA is only used in pre-finetuning. ‘VPA+PID’ uses standard PID and VPA, while ‘VPA+aPID’ combines VPA with adaptive PID. VPA+aPID shows the best learning efficiency, stability, and convergence.
@@ 4.3 Ablation Studies
We conduct experiments in various setups to better understand Marvel. As shown in Fig. 3, the performance is best when both VPA and aPID are used. In contrast, if only VPA is used with traditional dual ascent during online finetuning, it significantly slows down safe online learning and takes a much longer time to reduce the cost. If only aPID is applied without VPA, the learning performance is very similar to naive policy finetuning, which struggles to improve due to the erroneous Q-estimations. We also evaluate the effectiveness of adaptive control in aPID, by comparing the performance between Marvel (VPA+aPID) and Marvel with aPID replaced by PID (VPA+PID). In Fig. 3, the training curve for cost exhibits significant fluctuations without using aPID. More critically, when the cost is close to the limit, PID cannot reduce its control strength. As a result, even if on average the cost of VPA+PID is close to the threshold, it is very frequent that the real-time cost exceeds the limit substantially, which is in fact not safe. aPID can adjust $\lambda$ more precisely, which directly affects how the policy balances reward and cost at every timestep. In contrast, a delayed or inaccurate adjustment of the Lagrange multiplier alters the relative weighting of reward and cost, thereby degrading reward performance, as observed in the CarCircle and HalfCheetah plots (Appendix D.2).
Table 3: Performance comparison between VPA+aPID (which is standard Marvel), VPA, and Warm Start when agent have different performance of offline policy. Note that although the second line shows Marvel exceeding the cost threshold, the violation is minor and can be corrected with just a few update steps.
Environment VPA+aPID VPA Warm Start Offline
BallCircle 100 steps: r/c 580.3 / 19.2 579.0 / 18.4 54.4 / 0.0 166.0 / 11.0
500 steps: r/c 610.1 / 19.6 511.3 / 16.9 176.6 / 18.5
100 steps: r/c 543.6 / 20.5 576.8 / 92.1 430.7 / 13.5 446.0 / 6.0
500 steps: r/c 605.4 / 20.9 400.1 / 24.3 269.7 / 17.4
CarCircle 100 steps: r/c 283.8 / 20.0 320.1 / 49.7 107.3 / 8.8 145.0 / 17.0
500 steps: r/c 327.2 / 19.5 306.3 / 21.9 176.4 / 20.8
100 steps: r/c 330.6 / 18.8 260.6 / 19.6 73.2 / 71.1 265.0 / 14.0
500 steps: r/c 341.3 / 19.5 283.1 / 17.2 141.6 / 24.1
In practical O2O safe RL training, the quality of offline datasets can vary significantly, making it essential to evaluate whether Marvel maintains strong performance under different pretrained policies and Q-networks. As shown in Table 3, we train offline policies (and corresponding Q-networks) with varying quality levels. Marvel rapidly achieves near-optimal performance (with only a few online interaction steps, shown in Figure 9) regardless of the quality of the offline dataset or the effectiveness of the pretrained policy. This demonstrates the robustness of Marvel to variations in offline data quality. Furthermore, the algorithm consistently enforces cost constraints while improving rewards, thereby accommodating pretrained policies and Q-networks with diverse performance levels.
Regarding the necessity of each component of Marvel, Table 4 illustrates the importance of applying VPA to both the Q and Qc networks initialized from offline pre-trained models, as well as the contribution of the entropy term. Finetuning the pretrained Q-networks also outperforms learning new Q-networks from scratch in VPA, which implies that the pretrained Q-networks, although not accurate, can still provide meaningful prior knowledge.
The training curves corresponding to Table 3 and Table 4 are presented in Appendix D.2, which further validating Marvel’s robustness and design choices.: 1) Through extensive experiments across diverse environments, we confirm that the algorithm consistently maintains cost constraints while improving rewards across offline datasets of varying quality, which consequently accommodates offline pretrained policies and Q-networks with different performance levels. 2) The training dynamics confirm the rationality of VPA’s design, particularly the necessity of updating all Q-networks in conjunction with the entropy term. 3) Parameter sensitivity analyses in Appendix D.3 and Appendix D.4 reveal acceptable performance variance despite the introduced hyper-parameters. Furthermore, we verify correctness of our baseline implementation by demonstrating comparable performance to the standard library FSRL (Liu et al., 2024) under identical experimental settings, including interaction steps and policy update frequencies.
Offline Q Qc Entropy Pre-trained Q Performance
166.0 / 11.0 603.9 / 19.8
511.3 / 16.9
364.7 / 11.2
64.1 / 5.3
1.2 / 51.0
Table 4: Ablation study of VPA on BallCircle. “Offline” shows the initial policy/Q values. “Entropy” indicates whether entropy regularization is used. “Q”/“Qc” columns denote the targets of VPA. “Pre-trained Q” shows if VPA starts from an offline Q network. “Performance” reports final reward and cost.
@@ 5 Conclusion
O2O safe RL has great potential to put safe RL on the ground in real-world applications, by leveraging offline learning to facilitate fast online safe learning. In this paper, we proposed the first policy-finetuning based framework, namely Marvel, for O2O safe RL. In particular, by showing that naive finetuning would not work well, we identified two unique challenges in O2O safe RL, i.e., the erroneous Q-estimations and Lagrangian mismatch. To address these challenges, Marvel consisted of two key designs: 1) value pre-alignment to correct the Q-estimations before online finetuning, and 2) adaptive PID control to dynamically change the control parameters so as to rapidly and appropriately control the cost. Extensive experiments demonstrate the superiority of Marvel over multiple baselines. More importantly, Marvel is compatible to a variety of offline and online safe RL approaches, making it very practically appealing. We hope our work bridges the gap between offline and online safe RL, distinct from unconstrained RL, and enhances the efficiency of online safe RL, paving the way for practical applications.
Acknowledgement
We thank the anonymous reviewers and the Action Editor for their constructive feedback and helpful suggestions, which greatly improved the quality and clarity of this paper.
@@ Appendix A Overview of Algorithm
1:Input: Offline dataset $\mathcal{D}_{off}$, environment $\mathcal{E}$; offline RL loss functions $\{L_{off}^{Q_{\phi}},\,L_{off}^{Q_{c_{\phi_{c}}}},\,L_{off}^{\pi_{\theta}}\}$, and online RL loss functions $\{L_{on}^{Q_{\phi}},\,L_{on}^{Q_{c_{\phi_{c}}}},\,L_{on}^{\pi_{\theta}}\}$; offline learning rates $(\eta_{Q}^{off},\,\eta_{Q_{c}}^{off},\,\eta_{\pi}^{off})$; online learning rates $(\eta_{Q}^{on},\,\eta_{Q_{c}}^{on},\,\eta_{\pi}^{on})$; PID hyper-parameters and initial controller state $(\lambda_{0},I_{0},D_{0})$.
2:% These loss functions define a generic actor–critic safe RL algorithm: $L^{Q_{\phi}}$ and $L^{Q_{c_{\phi_{c}}}}$ are the objective functions for learning the reward critic $Q$ and the cost critic $Q_{c}$, respectively; $L^{\pi_{\theta}}$ is the policy optimization objective. Marvel treats them as modular components, allowing any safe RL algorithm to be plugged into the framework.
3:% In our implementation, we instantiate the offline losses using CPQ (Xu et al., 2022) and the online losses using SAC-lag (Ray et al., 2019), but Marvel is compatible with any actor–critic safe RL formulation.
4:Output: Trained policy parameters $\theta$, critic parameters $\phi,\phi_{c}$, online replay buffer $\mathcal{D}_{on}$, and final Lagrange multiplier $\lambda$.
5: Initialize network parameters $\phi,\phi_{c},\theta$; initialize $\mathcal{D}_{on}\leftarrow\emptyset$; set $(\lambda,I,D)\leftarrow(\lambda_{0},I_{0},D_{0})$.
6:while in offline training phase do
7:  Sample minibatch $(s,a,r,c,s^{\prime})\sim\mathcal{D}_{off}$
8:  $\phi\leftarrow\phi-\eta_{Q}^{off}\nabla_{\phi}L_{off}^{Q_{\phi}}$
9:  $\phi_{c}\leftarrow\phi_{c}-\eta_{Q_{c}}^{off}\nabla_{\phi_{c}}L_{off}^{Q_{c_{\phi_{c}}}}$
10:  $\theta\leftarrow\theta-\eta_{\pi}^{off}\nabla_{\theta}L_{off}^{\pi_{\theta}}$
11:end while
12:% VPA with offline dataset
13:while in VPA phase do
14:  for each VPA step do
15:   Sample transitions $(s,a,r,c,s^{\prime})\sim\mathcal{D}_{off}$
16:   Update $Q$ by Eq. 8, update $Q_{c}$ by Eq. 9
17:  end for
18:end while
19:while in online training phase do
20:  Roll out one episode in $\mathcal{E}$ with policy $\pi_{\theta}$, collect transitions $(s_{i},a_{i},r_{i},c_{i},s_{i+1})$ and append to $\mathcal{D}_{on}$
21:  Compute total episode cost $c_{t}$ and update EMA-smoothed cost $\tilde{c}_{t}$
22:  for each update step do
23:   Sample minibatch $(s,a,r,c,s^{\prime})\sim\mathcal{D}_{on}$
24:   $\phi\leftarrow\phi-\eta_{Q}^{on}\nabla_{\phi}L_{on}^{Q_{\phi}}$
25:   $\phi_{c}\leftarrow\phi_{c}-\eta_{Q_{c}}^{on}\nabla_{\phi_{c}}L_{on}^{Q_{c_{\phi_{c}}}}$
26:   $\theta\leftarrow\theta-\eta_{\pi}^{on}\nabla_{\theta}L_{on}^{\pi_{\theta}}$
27:   Update $\lambda$ by Eq. 12 with input $\tilde{c}_{t}$
28:   Update PID parameters by Eq. 15
29:  end for
30:end while